charges of +(10/3)*1/10000000 are placed at each of the


four corners of a square of side 8cm .the potential at the intersection of the diagonals is.


            PLEASE SOLVE THIS QUESTION.

Grade:12

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

 

Dear sunny shah,

V(r) = (1/4 π ε)*   (q1/r1+ q2/r2+ q3/r3+ q4/r4) = (1/4 π ε)*   4*(q/r) = (1/ π ε)*   (q/r)

 

q1=q2=q3=q4=q 

 

r=sqrt(4^2+4^2)   because  the total lenght of the side is 8cm . the cahrge is at diagonal of the square from the center

r=sqrt(32)  cm    chane to meter 

 

r=(1/10)*sqrt(32) m

ε= 8.854187817...×10−12 A·s/(V·m) = 8.854187817...×10−12 F/m

q=+(10/3)*1/10000000 = (10/3)*  10^-6

Π=22/7

 

now u can use the formula and values given above. do it so u can get practice and solve the otherparts

 


Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay

 

 


 

 

 

here